The decomposition of ethane (C2H6) to methyl radicals is a first-order reaction with a rate constant of 5.36 X 10^-4 s ^-1 at 700˚C: C2H6 ---> 2CH3. Calculate the half-life of the reaction in minutes.
Answer #1

We know that

for first order reaction

Half life t1/2 = 0.693 / rate constant k

So Half Life

t1/2 = 0.693 / (5.36 x 10-4 s-1)

= 1292.91 s

= 1292.91 /60 min

= 21.55 min

Therefore half life in min = 21.55 min